OpenStudy (anonymous):

For the polynomial below, 3 is a zero. H(x)=x^3-7x^2 +11x+3 Express h(x) as a product of linear factors.

OpenStudy (anonymous):

in otherwords, factor

OpenStudy (mathmale):

Hey, Bear: If 3 is a zero, then x-3 is a factor. There are several ways of factoring this polynomial. The long way is to do long division, that is, to divide (x-3) into x^3-7x^2+11x+3. There should be no remainder to this long division if x-3 actually is one of the factors. Another way, which I favor because it's so much faster, is synthetic division. Given that 3 is a zero, synth. div. should be a snap. Are you familiar with it? You identify the four coefficients of the x powers in the given polynomial: {1, -7, 11, 3}. ___________ Set up synthetic division as 3 | 1 -7 11 3 ___________ 1 Others or I will help if you encounter any difficulty with this synthetic division. Pls try it.

OpenStudy (anonymous):

Thanks. I'm not familiar with synthetic division but will give it a shot.

OpenStudy (mathmale):

Try long division first if that's familiar to you. Doing synthetic div after that will both demonstrate how you get the same numerical results with less work. _________________________ Long div.: x-3 | x^3 - 7x^2 +11x + 3

OpenStudy (anonymous):

x-3 is the denominator? I am watching Khan academy video in synthetic division and will see how it goes.

OpenStudy (anonymous):

I got \[1x ^{2}-4x-1+\frac{ 0 }{x-3 }\]

OpenStudy (campbell_st):

thats correct so h(x) = (x -3)(x^2 -4x -1) you need to find the linear factors of the quadaratic.. which will be complex conjugates

OpenStudy (campbell_st):

use the general quadratic formula on the quadratic factor \[x = \frac{4 \pm \sqrt{(-4)^2 - 4 \times 1 \times (-1)}}{2 \times 1}\] simplify this and then you'll have the other 2 linear factors.

OpenStudy (anonymous):

@campbell_st I simplified to x=\[-4\sqrt{3}\]

OpenStudy (campbell_st):

well I think you need to check it as its \[x = 2 \pm \sqrt{5}\] so the linear factors are \[h(x) = (x-3)(x - (2-\sqrt{5}))(x + (2 - \sqrt{5})\]